OpenStudy (anonymous):

Let \(T:\mathbb{R^n}\rightarrow\mathbb{R}^m\) and \(S:\mathbb{R^n}\rightarrow\mathbb{R}^m\) be linear. Show that \(S\circ T\) is also linear. Carefully justify each step in the process. I got this wrong and can't figure it out.

OpenStudy (loser66):

Let u1, u2 be arbitrary vectors in R^n and let c be a scalar, we must prove that S o T (u1+u2) = (SoT)(u1 ) + (SoT) (u2) and (SoT)(cu ) = c (SoT)(u) SoT(u1+u2 )= S (T (u1 + U2) = S (T(u1) + ST(u2) = SoT (u1) + SoT(u2) (close under addition) the part for scalar is too easy, right?

OpenStudy (zarkon):

the way you have it written \(S\circ T\) would not be defined

OpenStudy (zarkon):

double check the mappings of T and S

OpenStudy (anonymous):

You are correct, should be: \(T:\mathbb{R^n\rightarrow\mathbb{R}^m}\) and \(T:\mathbb{R^m\rightarrow\mathbb{R}^p}\).

OpenStudy (anonymous):

The second T should be S.

OpenStudy (zarkon):

that looks better ;)

OpenStudy (anonymous):

Thanks